### v1

# Ultimate Precision Limit of Quantum Thermal Machines

Abstract
Enhancing the precision of a thermodynamic process inevitably necessitates a thermodynamic cost–a principle formulated as the thermodynamic uncertainty relation. In stochastic thermodynamics, the thermodynamic uncertainty relation revealed a trade-off between the relative variance of a thermodynamic current and the entropy production. Although the thermodynamic uncertainty relation states that infinitely-large entropy production enables unbounded precision improvements, this is infeasible in reality. In this Letter, we establish the ultimate limits of precision for open quantum thermal machines operating within a finite-dimensional system and environment. We derive bounds on the relative variance and the expectation of observables, applicable to any unitary evolution of the composite system. These bounds are governed by the minimum eigenvalue factor, which serves as a maximal entropy production attainable by the quantum dynamics and parallels the concept of dynamical activity in classical stochastic thermodynamic. Our findings provide insights into fundamental limits on the precision of quantum thermal machines.
Introduction.—In the physical world, there is no free lunch. This fundamental principle is illustrated by the thermodynamic uncertainty relation, which shows that making a physical process more precise comes at a thermodynamic cost. It was first studied in classical stochastic thermodynamics [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] using Markov processes. Consider a classical Markov process, where $C$ represents a thermodynamic current over the time interval $[0,\tau]$, and $\mathbbl{\Sigma}$ denotes the entropy production dissipated during the interval. Then the following lower bound exists for the relative variance of $C$ [1, 2]:
$\frac{\mathrm{Var}[C]}{\mathbb{E}[C]^{2}}\geq\frac{2}{\mathbbl{\Sigma}}.$ (1)
where $\mathbb{E}[C]$ and $\mathrm{Var}[C]$ denote the expectation and variance of $C$, respectively. Equation (1) demonstrates that increasing the entropy production $\mathbbl{\Sigma}$ is necessary to improve the accuracy of a thermodynamic machine. Conversely, Eq. (1) indicates that if the entropy production could be increased infinitely, the accuracy of the thermal machine could be raised without limit. However, in reality, it is impossible to dissipate an infinite amount of entropy production due to the physical constraints. This issue becomes clearer when considering finite-dimensional quantum systems. The thermodynamic uncertainty relation of quantum thermodynamics has been a focus of active research in recent years [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Quantum thermodynamics [22, 23] aims to understand the fundamental principles governing the behavior of matter and energy at small scales. Open quantum dynamics can be described by the joint unitary evolution of the principal system $S$ and the environment $E$ with finite dimensions (Fig. 1(a)). Using the monotonicity of the quantum relative entropy and lower bounds on the divergences [24, 9, 25], it is not difficult to show (for more information on the settings, see Appendix A)
$\left(\frac{\sqrt{\mathrm{Var}_{f}[G]}+\sqrt{\mathrm{Var}_{i}[G]}}{\mathbb{E}_{f}[G]-\mathbb{E}_{i}[G]}\right)^{2}\geq\frac{1}{e^{\Sigma}-1}.$ (2)
Note that a similar bound for the quantum relative entropy (therefore the quantum entropy production) was recently derived in Ref. [26]. In Eq. (2), $\Sigma$ is the entropy production (cf. Eq. (23)) and $G$ is a measurement operator applied to the environment. $\mathbb{E}_{i}[G]$ and $\mathbb{E}_{f}[G]$ represent the expectations of $G$ with respect to the initial and final states, respectively. Similarly, $\mathrm{Var}_{f}[G]$ denotes the variance of $G$ in the final state, and $\mathrm{Var}_{i}[G]$ is defined analogously for the initial state. Equation (2) describing the thermodynamic uncertainty relation does not impose any limitations on the possible values of $\Sigma$ as long as it is non-negative; that is, achieving an infinitely precise thermal machine is formally possible if the entropy production becomes infinitely large. However, because the dimension of the composite system is finite, it is not difficult to imagine that an infinite entropy production is infeasible.
This Letter establishes the ultimate limit of precision for open quantum thermal machines operating within a finite-dimensional system and environment. We derive a lower bound for the expectation ratio of the norm of the observable (cf. Eq. (19)). This bound applies to any unitary operator acting on the composite system. Considering a specific case of the derived bound, we can obtain a lower bound for the relative variance (cf. Eq. (21)) that mirrors the form of the thermodynamic uncertainty relation. Additionally, we obtain an upper bound for the expectation of the observable (cf. Eq. (22)). These bounds are determined by the minimum eigenvalue factor. In addition to our main findings, we show that the minimum eigenvalue factor can be understood in two ways. First, it can be interpreted as the maximum entropy production for a given quantum system. Second, the minimum eigenvalue factor can be seen as analogous to the dynamical activity employed in classical Markov processes. These insights provide a deeper understanding of the minimum eigenvalue factor and its role in quantum systems.
Figure 1: (a) The joint unitary dynamics. The system $S$ and the environment $E$ undergoes the joint unitary operator $U$, in which the environment after the unitary is measured by the Hermitian operator $G$. (b) The energy spectrum of the system and the environment can be represented as two sets: $\{\sigma_{i}\}_{i=1}^{d_{S}}$ and $\{\epsilon_{i}\}_{i=1}^{d_{E}}$, where $d_{S}$ and $d_{E}$ denote the dimensions of the respective sets.
Results.—Let us consider an open quantum thermal machine the system $S$ and the environment $E$. We can consider the state transformation of the composite system by applying a joint unitary operator $U$. Let $H_{S}$ and $H_{E}$ be the Hamiltonian operators representing the system $S$ and the environment $E$, respectively. These operators can be expressed in their spectral decompositions as follows:
$H_{S}=\sum_{i=1}^{d_{S}}\sigma_{i}\ket{\sigma_{i}}\bra{\sigma_{i}},\quad H_{E}=\sum_{i=1}^{d_{E}}\epsilon_{i}\ket{\epsilon_{i}}\bra{\epsilon_{i}}.$ | . (3)
Here, $\sigma_{i}$ and $\epsilon_{i}$ denote the eigenvalues of $H_{S}$ and $H_{E}$ (Fig. 1(b)), respectively, while |$\ket{\sigma_{i}}$ and |$\ket{\epsilon_{i}}$ represent the corresponding eigenvectors. The dimensions of the system and the environment are given by $d_{S}$ and $d_{E}$, respectively. After preparing the initial states, we apply the unitary transformation $U$ to the composite system:
$\rho_{SE}^{\prime}=U(\rho_{S}\otimes\rho_{E})U^{\dagger},$ (4)
where $\rho_{S}$ and $\rho_{E}$ are the system and environmental initial states, respectively, and $\rho_{SE}^{\prime}$ is the joint state after the time evolution.
Let $\rho_{E}^{\prime}\equiv\mathrm{Tr}_{S}[\rho_{SE}^{\prime}]$ be the state of the environment after the unitary operation $U$. We wish to measure the environment after the unitary evolution (Fig. 1(a)). This approach is known as indirect measurement, where rather than directly measuring the system itself, an ancillary system is used to measure the system. By measuring the ancillary system, information about the principal system can be inferred. Continuous measurement [27], which is often considered in the context of the quantum thermodynamic uncertainty relation [17], falls into this category. Continuous measurements can be regarded as interactions between $S$ and $E$, where $S$ corresponds to the dynamics governed by the Lindblad equation, and $E$ corresponds to the field recording jump events. By measuring the field, information about the jumps can be obtained.
Let $G$ be an arbitrary observable that is to be applied to the environment $E$. Let $\mathbb{E}[G]\equiv\mathrm{Tr}[\rho_{E}^{\prime}G]$ be the expectation of $G$ and $\mathrm{Var}[G]\equiv\mathbb{E}[G^{2}]-\mathbb{E}[G]^{2}$ be the variance of $G$. Given the observable $G$, we want to obtain a lower bound of $\mathrm{Var}[G]/\mathbb{E}[G]^{2}$. Let $\lambda(G)$ be a set of eigenvalues of $G$, and $\lambda_{\min}(G)\equiv\min\lambda(G)$ and $\lambda_{\max}(G)\equiv\max\lambda(G)$. Although we aim to explore general observables, it is necessary to impose a condition on $G$; the minimum eigenvalue of $G$ is 00, $\lambda_{\min}(G)=0$. Without this condition, we can make the relative variance approach 00 by transforming $G$ to $G+a\mathbb{I}$ and letting $a$ go to infinity. This transformation allows the mean value of $G$, denoted as $\mathbb{E}[G]$, to become infinitely large while maintaining a constant variance, $\mathrm{Var}[G]$. Consequently, the relative variance $\mathrm{Var}[G]/\mathbb{E}[G]^{2}$ can be made infinitely small, which is undesirable for our analysis.
$\rho_{E}^{\prime}$ and $G$ admit the eigenvalue decomposition:
$\rho_{E}^{\prime}$ $=\sum_{n=1}^{d_{E}}r_{n}\ket{n}\bra{n},$ (5)
$G$ $=\sum_{n=1}^{d_{E}}g_{n}\ket{g_{n}}\bra{g_{n}},$ (6)
where $r_{n}\geq 0$ and $\sum_{n}r_{n}=1$. Here, we exclude the possibility of degeneracy for simplicity; however, this condition can be addressed if necessary. When a measurement is performed using the operator $G$, one of the eigenvalues $\lambda(G)=\{g_{1},g_{2},\ldots,g_{d_{E}}\}$ is observed. Since we have assumed that $\lambda_{\min}(G)=0$, we set $g_{1}=0$ without loss of generality. From the Petrov inequality [28], the lower bound to the relative fluctuation of $G$ is given by (see Appendix B)
$\frac{\mathbb{E}\left[G^{s}\right]^{r/(s-r)}}{\mathbb{E}\left[G^{r}\right]^{s/(s-r)}}\geq\frac{1}{1-P(G=0)},$ (7)
where $0<r<s$ and $P(G=0)\equiv\mathrm{Tr}[\rho_{E}^{\prime}\ket{g_{1}}\bra{g_{1}}]$ is the probability of observing $G=g_{1}=0$ given $\rho_{E}^{\prime}$. For $r=1$ and $s=2$, Eq. (7) can derive a lower bound for the relative variance:
$\frac{\mathrm{Var}[G]}{\mathbb{E}[G]^{2}}\geq\left(\frac{1}{P(G=0)}-1\right)^{-1}.$ (8)
Equations (7) and (8) suggest that minimizing $P(G=0)$ is essential for achieving higher precision. Since $P(G=0)$ depends on |$\ket{g_{n}}$, we should choose |$\ket{g_{n}}$ to minimize $P(G=0)$. We can express $P(G=0)=\sum_{n}r_{n}c_{n}$. Here, $c_{n}$ is defined by $c_{n}\equiv|\braket{n}{g_{1}}|^{2}$, and satisfies $c_{n}\geq 0$ and $\sum_{n}c_{n}=1$. Then the minimization problem can be formulated as follows:
$\underset{\{c_{n}\}}{\text{mimimize}}\quad$ $\sum_{n}r_{n}c_{n}$ (9)
subject to $1\displaystyle 0\leq c_{n}\leq 1,\sum_{n}c_{n}=$ 1 . (10)
This minimization problem can be solved with the linear programming. In linear programming, the optimal solution lies at an extreme point of the feasible region. In this case, extreme points are characterized by having exactly one of the $c_{n}$ equal to $1$ and the others equal to 00. The solution achieving the minimum is selected from these extreme points. Let $r_{1}$ be the minimum of $\lambda(\rho_{E}^{\prime})$, $r_{1}=\lambda_{\min}(\rho_{E}^{\prime})$. The the minimum of $P(G=0)$ is achieved when |$\ket{n=1}=\ket{g_{1}}$, which yields
$\frac{\mathbb{E}\left[G^{s}\right]^{r/(s-r)}}{\mathbb{E}\left[G^{r}\right]^{s/(s-r)}}\geq\frac{1}{1-r_{1}}=\frac{1}{1-\lambda_{\min}(\rho_{E}^{\prime})}.$ (11)
According to Eq. (11), minimizing the relative variance of the observable $G$ requires minimizing the smallest eigenvalue, denoted as $r_{1}$, of $\rho_{E}^{\prime}$. Note that we recently derived a thermodynamic uncertainty relation using the Petrov inequality, where the upper bound comprises the dynamical activity [29].
Next, we consider minimizing the minimal eigenvalue $r_{1}$ of $\rho_{E}^{\prime}.$ This problem setting is equivalent to the problem of thermodynamic cooling [30, 31, 32, 33, 34], where ancillary systems are used to cool the target system. $\rho_{E}^{\prime}$ is calculated by
$\rho_{E}^{\prime}=\sum_{n=1}^{d_{S}}\left(\bra{h_{n}}\otimes\mathbb{I}\right)\rho_{SE}^{\prime}\left(\ket{h_{n}}\otimes\mathbb{I}\right),$ (12)
where |$\ket{h_{n}}$ is a basis in $S$. Therefore, when considering all possible unitary transformations, the minimum achievable eigenvalue of $\rho_{E}^{\prime}$ is identical to the sum of the $d_{S}$ smallest elements of $\lambda(\rho_{SE}^{\prime})$ [30, 33]. Moreover, the eigenvalues of $\rho_{SE}^{\prime}$ is identical to the eigenvalues of $\rho_{SE}$, which is expressed by
$\lambda(\rho_{SE})=\{ab|a\in\lambda(\rho_{S}),b\in\lambda(\rho_{E})\}.$ (13)
[…]
$r_{1}=\lambda_{\min}(\rho_{E}^{\prime})\geq\sum_{i=1}^{d_{S}}\left[\lambda(\rho_{SE})\right]_{i}^{\uparrow}.$ (14)
Equation (14) states that a lower bound of the minimum eigenvalue can be expressed as the sum of the $d_{S}$ smallest eigenvalues of $\rho_{SE}$. Then the $d_{S}$ sum part in Eq. (14) can be bounded from below by
$\sum_{i=1}^{d_{S}}\left[\lambda(\rho_{SE})\right]_{i}^{\uparrow}$ $\geq d_{S}\lambda_{\min}(\rho_{SE})$
$=d_{S}\lambda_{\min}(\rho_{S})\lambda_{\min}(\rho_{E}).$ (15)
Combining Eq. (11) with (15), we obtain
$\frac{\mathbb{E}\left[G^{s}\right]^{r/(s-r)}}{\mathbb{E}\left[G^{r}\right]^{s/(s-r)}}\geq\frac{1}{1-d_{S}\lambda_{\min}(\rho_{S})\lambda_{\min}(\rho_{E})}.$ (16)
Figure 2: Results of numerical simulation. (a) Verification of Eq. (21). $\mathrm{Var}[G]/\mathbb{E}[G]^{2}$ is plotted against $(d_{E}e^{\Phi}/d_{S}-1)^{-1}$. The points represent random realizations, while the dashed line indicates the equality case of Eq. (21). (b) Verification of Eq. (22). $\mathbb{E}[G]$ is plotted against $\lambda_{\max}[G](1-d_{S}e^{\Phi}/d_{E})$. The points represent random realizations, while the dashed line indicates the equality case of Eq. (22). In (a) and (b), the system dimension is set to $d_{S}=2$. First, we randomly select the environment dimension $d_{E}$ from the set $\{2,3,4\}$. Then, we randomly generate the initial density matrix $\rho_{S}$ and the random Hamiltonian $H_{E}$. A randomly generated unitary $U$ is applied to the combined system. The measurement operator $G$ is randomly generated and apply the operation $G\to G-\lambda_{\min}[G]\mathbb{I}$ so that the minimum eigenvalue becomes 00.
Let $\gamma_{X}$ ($X\in\{S,E\}$) be the Gibbs state:
$\gamma_{X}=\frac{e^{-\beta H_{X}}}{Z_{X}(\beta)},\quad X\in\{S,E\},$ (17)
where $\beta$ is the inverse temperature and $Z_{X}(\beta)\equiv\mathrm{Tr}[e^{-\beta H_{X}}]$ ($X\in\{S,E\}$) is the partition function. Up to this point, no assumptions have been made regarding the initial state. From now on, the initial state of the environment will be taken as a Gibbs state, $\rho_{E}=\gamma_{E}$. This condition is often considered in open quantum systems [35]. In this scenario, $\lambda_{\min}(\gamma_{E})$ is bounded from below by [31]
$\lambda_{\min}\left(\gamma_{E}\right)=\frac{e^{-\beta\epsilon_{\max}}}{\mathrm{Tr}\left[e^{-\beta H_{E}}\right]}\geq\frac{e^{-\beta\epsilon_{\max}}}{d_{E}e^{-\beta\epsilon_{\min}}},$ (18)
where $\epsilon_{\max}\equiv\lambda_{\max}(H_{E})$ and $\epsilon_{\min}\equiv\lambda_{\min}(H_{E})$. Then we obtain
$\frac{\mathbb{E}\left[G^{s}\right]^{r/(s-r)}}{\mathbb{E}\left[G^{r}\right]^{s/(s-r)}}\geq\left(1-\frac{d_{S}}{d_{E}}e^{-\beta\Phi(\beta,\Delta\epsilon,\lambda_{\min}(\rho_{S}))}\right)^{-1},$ (19)
where $\Delta\epsilon\equiv\epsilon_{\max}-\epsilon_{\min}$ and we define
$\Phi(\beta,\Delta\epsilon,\lambda_{\min}(\rho_{S}))\equiv\ln\frac{1}{\lambda_{\min}(\rho_{S})}+\beta\Delta\epsilon.$ (20)
We refer to $\Phi$ as the minimum eigenvalue factor. For $s=2$ and $r=1$, Eq. (19) provides the lower bound to the relative variance:
$\frac{\mathrm{Var}[G]}{\mathbb{E}\left[G\right]^{2}}\geq\left(\frac{d_{E}}{d_{S}}e^{\Phi(\beta,\Delta\epsilon,\lambda_{\min}(\rho_{S}))}-1\right)^{-1},$ (21)
where $\Delta\epsilon\equiv\epsilon_{\max}-\epsilon_{\min}$ is the energy bandwidth of the environment $E$. Equation (21) shows that, to achieve a higher level of precision, the energy bandwidth $\Delta\epsilon$, quantified by the difference between the largest and smallest eigenvalues of the environment, should be large enough. A larger energy bandwidth results in greater energy dissipation; therefore, this aligns with our intuition. Later, we will demonstrate that the minimum eigenvalue factor is connected to the maximum entropy production. The lower bound in the conventional thermodynamic uncertainty relation comprises the entropy production, which depends on the actual time evolution of the system. However, the derived bound depends solely on the initial configuration and does not involve quantities related to the time evolution. Moreover, dimensions $d_{S}$ and $d_{E}$ play an important role in the bound. Moreover, taking $s=\infty$ and $r=1$, we have
$\mathbb{E}\left[G\right]\leq\lambda_{\max}[G]\left\{1-\frac{d_{S}}{d_{E}}e^{-\beta\Phi(\beta,\Delta\epsilon,\lambda_{\min}(\rho_{S}))}\right\}.$ (22)
Equation (22) states that the expectation of the observable $G$ relative to the maximum eigenvalue becomes larger for larger $\Phi$. This bound is reminiscent of a bound derived in Refs. [36, 29], where $\Phi$ should be replaced by the dynamical activity.
As mentioned in the introduction, a quantum thermodynamic uncertainty relation has a lower bound comprising the entropy production $\Sigma$. $\Phi$ as defined in Eq. (20) has the term $\beta\Delta\epsilon$. It is naturally expected that larger $\beta\Delta\epsilon$ also results in larger entropy production. Recently, Ref. [37] established a lower bound for the entropy production based on the minimum eigenvalue. We explore how the minimum eigenvalue serves as an upper bound for entropy production, offering a fundamental limit to thermodynamic precision. As before, suppose that the initial state of the environment is the Gibbs state $\rho_{E}=\gamma_{E}$. The quantum entropy production is defined by [38, 31]
$\Sigma=\mathcal{D}(\rho_{SE}(\tau)\|\rho_{S}(\tau)\otimes\gamma_{E}),$ (23)
where $\mathcal{D}(\rho\|\eta)$ is the quantum relative entropy:
$\mathcal{D}(\rho\|\eta)\equiv\mathrm{Tr}[\rho(\ln\rho-\ln\eta)].$ (24)
From Ref. [39], $\mathcal{D}(\rho\|\eta)$ is bounded from above by
$\mathcal{D}(\rho\|\eta)$ $\leq-S(\rho)-\ln\lambda_{\min}(\eta)$
$\leq-\ln\lambda_{\min}(\eta),$ (25)
where $S(\rho)\equiv-\mathrm{Tr}[\rho\ln\rho]$ is the von Neumann entropy. Using Eq. (25), $\Sigma$ is bounded from above by
$\Sigma$ $\leq-\ln\lambda_{\min}(\rho_{S}(\tau)\otimes\gamma_{E})$
$=-\ln\lambda_{\min}(\rho_{S}(\tau))-\ln\lambda_{\min}(\gamma_{E}).$ (26)
The minimum eigenvalue of $\gamma_{E}$ is given by Eq. (18). In the same way as in Eq. (15), a lower bound of $\lambda_{\min}(\rho(\tau))$ can be expressed by
$\lambda_{\min}(\rho_{S}(\tau))\geq d_{E}\lambda_{\min}(\rho_{S})\lambda_{\min}(\gamma_{E}).$ (27)
Therefore, the entropy production is bounded from above by
$\Sigma$ $\leq\ln\frac{d_{E}}{\lambda_{\min}(\rho_{S})}+2\beta\Delta\epsilon$
$=\ln d_{E}+\Phi(2\beta,\Delta\epsilon,\lambda_{\min}(\rho_{S})).$ (28)
Equation (28) shows that the minimum eigenvalue factor $\Phi$ is indeed related to the entropy production $\Sigma$; $\Phi$ provides an upper bound of $\Sigma-\ln d_{E}$ when we replace $\beta\to 2\beta$. By substituting the entropy production $\Sigma$ with the upper bound given in Eq. (28), we can derive a quantum thermodynamic uncertainty relation. However, the left side of Eq. (2) represents the relative variance of $G$ concerning the time-evolved state $\rho_{E}^{\prime}$ and the initial state $\rho_{E}$. This contrasts with Eq. (21), where the left-hand side is the relative variance at the final state.
Next, we show that $\Phi$ defined in Eq. (20) plays a similar role as the dynamical activity in classical Markov processes, which plays a central role in trade-off relations [3, 40, 5]. Let $W_{nm}$ be the transition rate from $m$th state to $n$th state in a classical Markov process and let $P_{m}(t)$ be the probability that the system is in $m$th state at time $t$. Given a Markov process $\dot{P}(t)=WP(t)$, where $W\equiv\{W_{nm}\}_{nm}$ and $P(t)=\{P_{m}(t)\}_{m}$ is the probability vector, the dynamical activity $\mathbbl{A}$ and the entropy production rate $\mathbbl{\Sigma}$ are defined by
$\mathbbl{\Sigma}$ $\equiv\sum_{n<m}\left(P_{m}W_{nm}-P_{n}W_{mn}\right)\ln\frac{P_{m}W_{nm}}{P_{n}W_{mn}},$ (29)
$\mathbbl{A}$ $\equiv\sum_{n<m}\left(P_{m}W_{nm}+P_{n}W_{mn}\right).$ (30)
It has been revealed that the dynamical activity constitute an upper bound to the entropy production in the steady-state condition [41]:
$\mathbbl{\Sigma}\leq\kappa(R)\mathbbl{A},$ (31)
where $R=\max_{n\neq m}W_{mn}/W_{nm}$ and $\kappa(R)\equiv(\ln R)(R-1)/(R+1)$. The entropy production can become arbitrarily large if we allow the jump rate ratio $W_{mn}/W_{nm}$ to be infinitely large. Equation (31) states that, when there is an upper bound to the ratio $W_{mn}/W_{nm}$, as specified by $R$, the entropy production $\mathbbl{\Sigma}$ is bounded from above by the dynamical activity $\mathbbl{A}$. References [29, 42] showed that $\mathbbl{A}$ constitutes a lower bound in a classical thermodynamic uncertainty relation. Let $C$ be an arbitrary trajectory observable within the time interval $[0,\tau]$, which should vanish when there is no jump. Then the following relation is known to hold [29, 42]:
$\frac{\mathrm{Var}[C]}{\mathbb{E}\left[C\right]^{2}}\geq\left(e^{\mathbbl{A}\tau}-1\right)^{-1}.$ (32)
Equation (32) shows that the system provides more precise measurements when $\mathbbl{A}$ is larger. Comparison of Eqs. (21) and (32) and comparison of Eqs. (28) and (31) suggest that $\Phi$, as defined in Eq. (20), plays a role similar to that of the dynamical activity. Specifically, both $\Phi$ and $\mathbbl{A}$ serve as lower bounds in trade-off relations and as upper bounds for entropy production.
The bound shown in Eq. (21) is significant from a thermodynamic perspective, particularly in relation to the third law of thermodynamics. In recent years, there has been growing interest in verifying theoretical thermodynamic relations using real quantum computers [43, 44, 45]. In our previous study [46], we confirmed a variant of the quantum thermodynamic uncertainty relation using the IBM quantum computer. However, due to the third law of thermodynamics, it is impossible to prepare a pure state as the initial state [30, 31, 33, 47, 47]. As a result, the initial state becomes mixed and thus $\lambda_{\min}(\gamma_{E})>0$, which implies that $\beta<\infty$. Consequently, even when considering the third law of thermodynamics, Eq. (21) may represent a possible limit on the achievable accuracy.
Numerical simulation.—We conduct numerical simulations to verify Eq. (21). To do this, we randomly generate the system density matrix $\rho_{S}$ and the environmental Hamiltonian $H_{E}$ for an initial Gibbs state $\rho_{E}$. We then create a random unitary $U$ and a random Hermitian operator $G$, whose minimum eigenvalue is 0. Repetition of this process multiple times, we obtain random realizations for both sides of Eq. (21): the left-hand side, $\mathrm{Var}[G]/\mathbb{E}[G]^{2}$, and the right-hand side, $\left(d_{E}e^{\Phi}/d_{S}-1\right)^{-1}$, which are plotted as points in Fig. 2(a). The dashed line in Fig. 2(a) represents the equality condition of the bound. As we can see, all the points lie above this line, indicating that Eq. (21) holds true in our simulations. Similarly, in Fig. 2(b), we verify Eq. (22) by plotting the left-hand side, $\mathbb{E}[G]$, against the right-hand side, $\lambda_{\max}[G]\left(1-d_{S}e^{-\Phi}/d_{E}\right)$. Once again, the dashed line in Fig. 2(b) represents the equality case of Eq. (16), and our numerical results confirm the validity of Eq. (22).
Conclusion.—In this Letter, we established fundamental precision limits for quantum thermal machines operating in the finite-dimensional system and environment. We derived a universal lower bound for the relative variance and the expectation, and introduced the minimum eigenvalue factor $\Phi$, which determines these bounds and serves as an analog to the dynamical activity in classical Markov processes. Our findings reveal that, unlike classical systems, where infinite precision might be achievable with infinite entropy production, quantum systems face fundamental limitations due to their finite-dimensional nature. Our study provides a comprehensive framework for understanding the ultimate precision limits in quantum thermal machines.
Appendix A Quantum relative entropy thermodynamic uncertainty relation
Quantum thermodynamic uncertainty relation based on the quantum relative entropy was considered in Ref. [26]. We here show a variant using the monotonicity and the variance-mean lower bound. For arbitrary density operators $\rho$ and $\eta$, let $F(\rho,\eta)$ be the quantum fidelity:
$F(\rho,\eta)\equiv\left(\mathrm{Tr}\sqrt{\sqrt{\rho}\eta\sqrt{\rho}}\right)^{2}.$ (33)
According to Theorem 7 in Ref. [48], the following relation holds between $F(\rho,\eta)$ and $\mathcal{D}(\rho||\eta)$:
$e^{-D(\rho\|\eta)}\leq F(\rho,\eta).$ (34)
We next derive an upper bound of the quantum fidelity using the probabilities associated with the measurement $G$ on the environment $E$. Upon conducting a measurement with the operator $G$, one of the eigenvalues from the set $\lambda(G)=\{g_{1},g_{2},\ldots,g_{d_{E}}\}$ is observed. Let $\Pi(g_{n})$ and $P(g_{n})$ be the probability of measuring $g_{n}$ with $\rho_{E}$ and $\rho_{E}^{\prime}$, respectively, that is $\Pi(g_{n})=\mathrm{Tr}[\rho_{E}\ket{g_{n}}\bra{g_{n}}]$ and $P(g_{n})=\mathrm{Tr}[\rho_{E}^{\prime}\ket{g_{n}}\bra{g_{n}}]$. Moreover, the fidelity satisfies the following relation:
$\sqrt{F(\rho_{E}^{\prime},\rho_{E})}$ $\leq\sum_{n}\sqrt{P(g_{n})\Pi(g_{n})}$
$=1-\mathrm{Hel}^{2}\left(P,\Pi\right).$ (35)
where $\mathrm{Hel}^{2}(\cdot,\cdot)$ is the Hellinger distance:
$\mathrm{Hel}^{2}\left(P,\Pi\right)\equiv\frac{1}{2}\sum_{n}\left(\sqrt{P(g_{n})}-\sqrt{\Pi(g_{n})}\right)^{2}.$ (36)
Recall that $\Sigma=\mathcal{D}(\rho_{SE}(\tau)\|\rho_{S}(\tau)\otimes\rho_{E})\geq\mathcal{D}(\rho_{E}^{\prime}||\rho_{E})$ due to the monotonicity for $\rho_{E}=\gamma_{E}$. Using the lower bound of the Hellinger distance proved in Ref. [25], we obtain
$\mathrm{Hel}^{2}\left(P,\Pi\right)\geq 1-\left[\left(\frac{\mathbb{E}_{f}[G]-\mathbb{E}_{i}[G]}{\sqrt{\mathrm{Var}_{f}[G]}+\sqrt{\mathrm{Var}_{i}[G]}}\right)^{2}+1\right]^{-\frac{1}{2}},$ (37)
where
$\mathrm{E}_{f}[G]\equiv\sum_{n}g_{n}P(g_{n}),\mathbb{E}_{i}[G]\equiv\sum_{n}g_{n}\Pi(g_{n}).$ (38)
The variances $\mathrm{Var}_{f}[G]$ and $\mathrm{Var}_{i}[G]$ are defined in a similar way. Using Eqs. (34), (35) and (37), the following relation holds:
$e^{-\mathcal{D}(\rho_{E}^{\prime}\|\rho_{E})}$ $\leq F(\rho_{E}^{\prime},\rho_{E})$
$\leq\left[\left(\frac{\mathbb{E}_{f}[G]+\mathbb{E}_{i}[G]}{\sqrt{\mathrm{Var}_{f}[G]}-\sqrt{\mathrm{Var}_{i}[G]}}\right)^{2}+1\right]^{-1},$ (39)
which directly yields Eq. (2).
Appendix B Petrov inequality
Let $X$ be a random variable. The Petrov inequality [28] is given by
$P(|X|>b)\geq\frac{\left(\mathbb{E}\left[|X|^{r}\right]-b^{r}\right)^{s/(s-r)}}{\mathbb{E}\left[|X|^{s}\right]^{r/(s-r)}},$ (40)
where $0<r<s$, $b\geq 0$, and the condition $b^{r}\leq\mathbb{E}\left[|X|^{r}\right]$ must be satisfied. By setting $r=1$, $s=2$, and $b=\lambda\mathbb{E}[|X|]$ in this inequality, where $0\leq\lambda\leq 1$, we can derive the well-known Paley-Zygmund inequality [49]. By definition, the observable $G$ is non-negative (see Eq. (6) and the following assumption). Therefore, By setting $b=0$ in Eq. (40), we have
$\frac{\mathbb{E}\left[G^{s}\right]^{r/(s-r)}}{\mathbb{E}\left[G^{r}\right]^{s/(s-r)}}\geq\frac{1}{1-P(G=0)}.$ (41)
When taking $r=1$ and $s=2$, we have
$\frac{\mathrm{Var}\left[G\right]}{\mathbb{E}\left[G\right]^{2}}\geq\frac{P(G=0)}{1-P(G=0)}.$ (42)
When taking $r=1$ and $s\to\infty$, we have
$\left(1-P(G=0)\right)\lambda_{\max}(G)\geq\mathbb{E}\left[G\right].$ (43)
Here, we used $\mathbb{E}[G^{s}]^{1/s}\overset{s\to\infty}{\longrightarrow}\lambda_{\max}(G)$.
Appendix C Different initial settings
Figure 3: The joint unitary dynamics. The system $S$ and the environment $E$ undergoes the joint unitary operator $U$, in which the environment after the unitary is measured by the Hermitian operator $G$. The initial states are (a) $\rho_{S}=\gamma_{S}$ and (b) $\rho_{S}=\gamma_{S}$ and $\rho_{E}=\gamma_{E}$.
In the main text, the initial state of the environment was assumed to be a Gibbs state. It is possible to consider cases where the main system is in a Gibbs state or where both the main system and the environment are in Gibbs states.
Consider the case where the initial state of the main system $S$ is the Gibbs state, $\rho_{S}=\gamma_{S}$ (Fig. 3(a)). In this case, $\lambda_{\min}(\gamma_{S})$ can be bounded from below by
$\lambda_{\min}\left(\gamma_{S}\right)=\frac{e^{-\beta\sigma_{\max}}}{\mathrm{Tr}\left[e^{-\beta H_{S}}\right]}\geq\frac{e^{-\beta\sigma_{\max}}}{d_{S}e^{-\beta\sigma_{\min}}}.$ (44)
By using Eq. (44), we can repeat the same argument. Using the Petrov inequality, we have
$\frac{\mathbb{E}\left[G^{s}\right]^{r/(s-r)}}{\mathbb{E}\left[G^{r}\right]^{s/(s-r)}}\geq\frac{1}{1-e^{-\Psi(\beta,\Delta\sigma,\lambda_{\min}(\rho_{E}))}},$ (45)
where
$\Psi(\beta,\Delta\sigma,\lambda_{\min}(\rho_{E}))\equiv-\ln\lambda_{\min}(\rho_{E})+\beta\Delta\sigma.$ (46)
Equation (46) corresponds to $\Phi$ defined in Eq. (20). Again, the energy bandwidth $\Delta\sigma\equiv\sigma_{\max}-\sigma_{\min}$ of the system Hamiltonian $H_{S}$ plays an important role, but unlike the case of Eq. (19), the dimensions of the system, $d_{S}$ and $d_{E}$, are not included explicitly. Setting $r=1$ and $s=2$, the bound for the relative variance is obtained:
$\frac{\mathrm{Var}\left[G\right]}{\mathbb{E}\left[G\right]^{2}}\geq\left(e^{\Psi(\beta,\Delta\sigma,\lambda_{\min}(\rho_{E}))}-1\right)^{-1}.$ (47)
where $\Delta\sigma\equiv\sigma_{\max}-\sigma_{\min}$. Again using $s=\infty$ and $r=1$, we obtain the expectation bound:
$\mathbb{E}[G]\leq\lambda_{\max}[G]\left(1-e^{-\Psi(\beta,\Delta\sigma,\lambda_{\min}(\rho_{E}))}\right).$ (48)
Next, we consider the scenario where both the initial states of the system and the environment are Gibbs states, denoted as $\rho_{S}=\gamma_{S}$ and $\rho_{E}=\gamma_{E}$ (Fig. 3(b)). Then we obtain
$\frac{\mathbb{E}\left[G^{s}\right]^{r/(s-r)}}{\mathbb{E}\left[G^{r}\right]^{s/(s-r)}}\geq\frac{1}{1-e^{-\Omega(d_{E},\Delta\sigma,\Delta\epsilon)}},$ (49)
where $\Omega(d_{E},\Delta\sigma,\Delta\epsilon)$ is defined by
$\Omega(d_{E},\Delta\sigma,\Delta\epsilon)=\ln d_{E}+\beta\left(\Delta\sigma+\Delta\epsilon\right).$ (50)
Again, by setting $r=1$ and $s=2$, we obtain
$\frac{\mathrm{Var}\left[G\right]}{\mathbb{E}\left[G\right]^{2}}\geq\frac{1}{e^{\Omega(d_{E},\Delta\sigma,\Delta\epsilon)}-1}.$ (51)
By setting $r=1$ and $s=\infty$, we obtain
$\mathbb{E}[G]\leq\lambda_{\max}[G]\left(1-e^{-\Omega(d_{E},\Delta\sigma,\Delta\epsilon)}\right).$ (52)
Acknowledgements.
This work was supported by JSPS KAKENHI Grant Number JP23K24915.